OpenStudy (anonymous):

Solve 2|5x − 2| + 4 = 18 for x. Show your work to receive full credit. (absolute value)

OpenStudy (anonymous):

|10x - 4| + 4 =18 |10x - 4| = 14 10x= 14 + 4 10x=18 x=1.8

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

no problem ;)